# Contrastive Visual Data Augmentation

### v1

@@ Abstract
Large multimodal models (LMMs) often struggle to recognize novel concepts, as they rely on pre-trained knowledge and have limited ability to capture subtle visual details. Domain-specific knowledge gaps in training also make them prone to confusing visually similar, commonly misrepresented, or low-resource concepts. To help LMMs better align nuanced visual features with language, improving their ability to recognize and reason about novel or rare concepts, we propose a Contrastive visual Data Augmentation (CoDA) strategy. CoDA extracts key contrastive textual and visual features of target concepts against the known concepts they are misrecognized as, and then uses multimodal generative models to produce targeted synthetic data. Automatic filtering of extracted features and augmented images is implemented to guarantee their quality, as verified by human annotators. We show the effectiveness and efficiency of CoDA on low-resource concept and diverse scene recognition datasets including INaturalist and SUN. We additionally collect NovelSpecies, a benchmark dataset consisting of newly discovered animal species that are guaranteed to be unseen by LMMs. LLaVA-1.6 1-shot updating results on these three datasets show CoDA significantly improves SOTA visual data augmentation strategies by 12.3% (NovelSpecies), 5.1% (SUN), and 6.0% (iNat) absolute gains in accuracy.  [^1] Code and data will be released upon publication acceptance.
@@ 1 Introduction
Recent advancements in multimodal pre-training (OpenAI, 2023; Google, 2023; Hurst et al., 2024) and visual instruction tuning (Liu et al., 2023b, a, 2024c) have enabled impressive LMM abilities. However, as shown in Fig.1, it still remains a challenge for current state-of-the-art proprietary and open-source models to robustly recognize novel visual concepts (e.g.“Clouded Tiger Cat” Fig.1a) and confusing / low-resource / commonly misrepresented visual concepts (e.g. “Resupply Base” Fig.1b).
In order to help models better acquire new visual concepts and distinguish confusable concepts, existing approaches straightforwardly 1). Fine-tune text decoder on new textual corpora to expand the concept base; and 2). Fine-tune both vision and text components on new web image-text pairs for visual concept acquisition. These approaches are ineffective due to scarcity of data for certain concepts and inefficiency caused by not knowing what precisely confused the models (3.1). As depicted in Fig.1, it can be difficult to obtain ample high-quality real images for novel concepts such as new animal species. While for confusing concepts, the problem usually lies with biased concept representation in web image-text data. For example: online images of “Resupply Base” mostly only consist of exterior views of the architecture without the interior details, which may cause the models to confuse it with a “Wholesale Store” that shares some interior features.
To help LMMs recognize and reason about novel and confusing concepts more robustly and efficiently, we propose CoDA, a Contrastive Visual Data Augmentation technique. For each target concept, CoDA first identifies a “confusable concept” that the LMM finds most similar to the target. Then, it extracts contrastive textual and visual features of the target concept with respect to the confusable concept. The extracted features go through a filtering process based on discriminability and generability to make sure that: 1). The features are possessed by the target concept but not the confusable concept; and 2). The feature can be reliably generated by the text-to-image generative model and recognized by the LMM. Afterwards, the features are passed to the text-to-image generative model to produce augmented visual instances of the target concept. To make sure that the features are indeed generated and recognizable by the LMM, CoDA again uses the LMM’s zero-shot inference to rank and filter the augmented images. Finally, the resulting augmented images can be used to update the LMM via low-rank adaptation, basic fine-tuning, in-context learning, or any other method of choice.
@@
Comprehensive experiments with LLaVA-NeXT on the 3 datasets show CoDA performs surprisingly well in teaching LMMs novel and confusing concepts, significantly improving data and compute efficiency compared to existing methods. In additional experiments, we show that CoDA is also able to improve novel concept recognition for traditional classifiers like ViT and proprietary LMMs such as GPT4o-mini. Finally, ablation experiments show that CoDA can be significantly improved by simply replacing its off-the-shelf components such as the text-to-image generation model with superior versions of similar models.
@@
  • NovelSpecies, a new benchmark dataset of novel animal species discovered in recent years, providing an ideal benchmark for novel concept recognition. NovelSpecies currently consisting of 2240 annotated images and will continue to be updated with future discoveries.
@@ 2 Related Works
Few-shot image recognition is a long-standing problem in the vision community. Early works in this area focused on improving traditional image classifiers on classifying existing concepts (Vinyals et al., 2016; Finn et al., 2017; Nichol, 2018; Dhillon et al., 2019; Tian et al., 2020; Bhagat et al., 2023; Afrasiyabi et al., 2022). On the other hand, while recent advancements in the training of vision language models (VLMs) and large multimodal models (LMMs) (OpenAI, 2023; Google, 2023; Hurst et al., 2024; Liu et al., 2023b, a, 2024c) have shown great promise and extensibility, they still severely lag behind traditional models in image classification, especially for low-resource, novel, and confusing concepts (Zhang et al., 2024; Cooper et al., 2024).
While commonly used text-side VLM data augmentation strategies (Yuksekgonul et al., 2022; Yang et al., 2023; Liu et al., 2024e; Sharifzadeh et al., 2024) have little effect on this issue, a more promising technique to solve this is through visual data augmentation. This includes basic visual manipulations such as cropping, flipping, and rotation (Yang et al., 2022; Kumar et al., 2024); and more advanced model-based augmentation such as style transfer (Zheng et al., 2019; Chun & Park, 2021) and image mixing (Uddin et al., 2020; Xie et al., 2021; Hao et al., 2023). More recently, with the rise of controllable and promptable visual generative models, knowledge and feature editing-based augmentation methods (Liu et al., 2022; Wu et al., 2023; Jin et al., 2024) have gained in popularity. Such methods generally focus on using multimodal data and general knowledge bases to guide image-editing models in creating augmented visual data based on existing images.
@@ 3 Methods
As shown in Fig.2, CoDA consists of 4 major steps including contrastive textual and visual feature extraction, feature filtering, feature-controlled image generation, and augmented image filtering. Together these steps ensure CoDA reliably generates informative and high-quality augmented images that help LMMs recognize novel and confusing concepts.
@@ 3.1 Feature Extraction
In our exploratory experiments, we find that significant mis-recognition errors occur on low-resource or commonly mis-represented concepts in vision-language instruction fine-tuning and multimodal pre-training datasets, which the LMMs are trained on. For example, the LLaVA 1.6 (34B) model (Liu et al., 2024c), mainly tuned on LAION-GPT-4V(LAION, 2024) and ShareGPT-4V (Chen et al., 2023) datasets, has a strong tendency to mis-recognize interior images of “Resupply Base” as “Wholesale Store” (Fig.1). Unsurprisingly, we find that all related references of “Resupply Base” across the 3 instruction-tuning datasets only depict exterior views of the concept rather than interior views. While the concept itself is not a low-resource concept in existing text corpora, it is severely low-resource and also commonly mis-represented in vision-language instruction fine-tuning datasets.
@@
While textual feature extraction generally works well for pre-existing and non-hyper-domain-specific concepts that are prevalent in textual data sources, it tends to fail when either of the conditions are not met. For example, a large language model with a knowledge cutoff prior to June 2023 would not be able to provide meaningful features regarding the Apple Vision Pro device announced in July, or the new animal species “Clouded Tiger Cat (L. pardinoides)” first described by scientists in April 2024 (1). In addition to this weakness, LLM-based textual feature extraction is also unreliable when asked to provide detailed information regarding hyper-domain-specific concepts like the “Mazda MX-5 Miata RF” or the “Lear’s Macaw (Anodorhynchus Leari)”. In practice, we observe that for novel and hyper-domain-specific concepts, most of the LLM extracted textual features end up being filtered out by our automatic feature filtering module.
@@ 3.2 Feature Filtering
    Discriminability ($D(f,\mathcal{C}_{T},\mathcal{C}_{M})$) : measures whether a feature $f$ can indeed be used to differentiate the target class $\mathcal{C_{\mathcal{T}}}$ from the misidentified concept $\mathcal{C_{\mathcal{M}}}$ ($f$ must first be a valid feature of $\mathcal{C_{\mathcal{T}}}$).
@@
Here we use an equal number of images of the target and misidentified concepts. A score below 0.5 indicates that the feature is more likely to be associated with the misidentified class rather then the target class. To ensure that selected features are more strongly associated with the target class, we filter out all features with Discriminability below 0.6. This method avoids the CLIP score bias against smaller features by only comparing feature association with the two classes and not relying on the absolute CLIP score.
Generability is calculated in a similar manner, comparing the average CLIP similarity between $f$ and synthetic images of the target concept against the average CLIP similarity between $f$ and real images of the misidentified concept:
$G(f,\mathcal{C}_{T},\mathcal{C}_{M})=\sum_{i\in I}\frac{\text{CLIP}(f,i_{\mathcal{C}_{T}}^{\text{synthetic}})}{\text{CLIP}(f,i_{\mathcal{C}_{T}}^{\text{synthetic}})+\text{CLIP}(f,i_{\mathcal{C}_{M}}^{\text{real}})}$
@@ 3.3 Image Generation and Verification
To ensure final images for augmentation contain our extracted and filtered target concept features, we propose a simple automatic verification metric that checks whether desired features are recognized in the augmented images by the LMM we want to update: Given the vanilla LMM $\mathcal{M}$, a set of features $\mathcal{F}$, and an augmented images $i^{\text{synthetic}}$, the feature satisfaction rate $S(i^{\text{synthetic}},F,M)$ for each augmented image:
@@ 3.4 Human Evaluation
Results in Tab.1 show human annotators overwhelmingly agree that the final extracted features belong to the target concept (92%) but not the misidentified concept (14%). The augmented synthetic images of the target concept also likely contains the desired features (83%), though as expected, there is some information loss between the text-to-image generation step. In addition, the three independent annotators generally agreed in their response (>0.8 IAA).
@@ 4 NovelSpecies Dataset
To create NovelSpecies, we start by collecting the list of species first described in each year by Wikidata (Wikidata, 2024). Then, to make sure we can curate a visual benchmark of novel species, we manually annotate and filter out extinct species and species with too few publicly available images. After filtering, we end up with a dataset of 64 new species, each consisting of 35 human-verified image instances, thus a total of 2240 images. The images are split into training, validation, and test sets. For each specie, there are 5 training images, 15 validation images, and 15 test images. This data split is consistent with our goal of creating a benchmark dataset for novel concept recognition, where the maximum number of training instances for a completely unseen concept can range from 1 to 5.
@@ 5.1 Datasets and Baselines
In additional to these 3 baselines, we also include ablations of CoDA with non-contrastive textual and visual features, i.e. w/o contrastive guidance from confusable concepts (3.1) nor discriminability-based feature filtering (3.2).
@@ 5.2 Main Experiment
Experiment results across the 3 datasets show that CoDA consistently outperforms existing traditional and feature-based data augmentation methods in the Fixed Real Data setting. When augmenting the training set with just a single synthetic image, CoDA is able to achieve 11.8% (NovelSpecies), 10.0% (SUN), and 17.8% (iNat) absolute gains in accuracy compared using all real images. It further out-performs the best existing baseline augmentation methods by 5-12% absolute gains. We also observe that ablated performance of CoDA (w/o contrastive) is still significantly above traditional and image-editing-based augmentation baselines while being almost consistently below CoDA’s performance. This shows the benefits of text-to-image generative augmentation methods compared to existing methods, as well as the benefits of fine-grained textual features during inference. This also highlights the need for contrastive feature selection and discriminability-based feature filtering. We find that increasing the number of augmented synthetic images does not necessarily improve updated model performance, this may be attributed to the fact that all generated images are ranked and selected from the same pool, with the first image being of the highest quality. Finally, the largest improvement over existing baselines can be seen in NovelSpecies, where CoDA methods involving visual features achieve the highest performance. This makes sense as the visual feature extraction method is designed to be robust to novel concepts with little textual documentation.
@@ 5.3 Additional Experiments
As explained in Sec.3, off-the-shelf model components used in CoDA can be easily swapped for superior versions of similar models to improve performance. To demonstrate this, we replace the open-weight Stable Diffusion 3.5 Large Turbo model (Stability AI, 2024) with the SOTA proprietary Recraft V3 Model (Recraft.AI, 2024) and run the same LLaVA-updating experiments as in Tab.2. Here we note that Recraft V3 has better instruction-following ability as well as better image generation quality compared to Stable Diffusion 3.5 Large Turbo. More details on these differences can be found in Sec.A.1. Our experiment results in Tab.3 show a significant performance boost when LoRA fine-tuning LLaVA with Recraft V3 produced synthetic images compared to fine-tuning on all-real data (28.7%) and also compared to fine-tuning on Stable Diffusion 3.5 Large Turbo produced synthetic data (7.9%). This demonstrates the potential increase of CoDA’s effectiveness along with improvements in Text-to-Image generative models. We believe it is also possible to achieve similar improvements by replacing the LLM/VLM components of CoDA with superior models in the future.
@@
While proprietary LMMs like GPT4o-mini (Hurst et al., 2024) tend to have relatively strong 0-shot performance on existing datasets such as SUN and iNaturalist, their performance significantly degrades on NovelSpecies due to having never encountered the novel concepts. To test whether CoDA can effectively improve novel concept recognition performance for such Proprietary LMMs, we fine-tune the gpt-4o-mini-2024-07-18 model using CoDA and relevant augmentation baselines. Results in Tab.3 demonstrate a significant performance gain (9.5%) for GPT4o-mini after being fine-tuned on CoDA augmented synthetic images. While this improvement is not as significant compared to the LLaVA-1.6 model (20.3%), it is due to GPT4o-mini’s better base performance.
@@
In addition to evaluating CoDA on LMMs which take image-text input and produces text output, we also test whether it can help traditional image classifiers recognize novel concepts. We run the widely-used ViT classifier (Alexey, 2020) on NovelSpecies with CoDA and other augmentation baselines. Results in Tab.3 show that CoDA is able to achieve a consistent performance gain over existing baselines for ViT-base (9.1% for single-shot augmentation). The ViT classifier provides stronger base performance compared to general VLMs, thus offering less room for improvement. However, we note here that our main focusing on improving LMMs instead of such traditional classifiers stems from LMMs’ superior extensibility and generalizability to other related tasks such as recognition-based reasoning and explanation.

### v2

@@ Abstract
Large multimodal models (LMMs) often struggle to recognize novel concepts, as they rely on pre-trained knowledge and have limited ability to capture subtle visual details. Domain-specific knowledge gaps in training also make them prone to confusing visually similar, commonly misrepresented, or low-resource concepts. To help LMMs better align nuanced visual features with language, improving their ability to recognize and reason about novel or rare concepts, we propose a Contrastive visual Data Augmentation (CoDA) strategy. CoDA extracts key contrastive textual and visual features of target concepts against the known concepts they are misrecognized as, and then uses multimodal generative models to produce targeted synthetic data. Automatic filtering of extracted features and augmented images is implemented to guarantee their quality, as verified by human annotators. We show the effectiveness of CoDA on low-resource concept and diverse scene recognition datasets including INaturalist and SUN. We additionally collect NovelSpecies, a benchmark dataset consisting of newly discovered animal species that are guaranteed to be unseen by LMMs. LLaVA-1.6 1-shot updating results on these three datasets show CoDA significantly improves SOTA visual data augmentation strategies by 12.3% (NovelSpecies), 5.1% (SUN), and 6.0% (iNat) absolute gains in accuracy. Code and data at contrastive-visual-data-augmentation.github.io
@@ 1 Introduction
Recent advancements in multimodal pre-training (OpenAI, 2023; Google, 2023; Hurst et al., 2024) and visual instruction tuning (Liu et al., 2023b, a, 2024c) have enabled impressive LMM abilities. However, as shown in Figure 1, it still remains a challenge for current state-of-the-art proprietary and open-source models to robustly recognize novel visual concepts (e.g.“Clouded Tiger Cat” Figure 1a) and confusing / low-resource / commonly misrepresented visual concepts (e.g. “Resupply Base” Figure 1b).
In order to help models better acquire new visual concepts and distinguish confusable concepts, existing approaches straightforwardly 1). Fine-tune text decoder on new textual corpora to expand the concept base; and 2). Fine-tune both vision and text components on new web image-text pairs for visual concept acquisition. These approaches are ineffective due to data scarcity for certain concepts and data inefficiency caused by not knowing what precisely confused the models (3.1). As depicted in Figure 1, it can be difficult to obtain ample high-quality real images for novel concepts such as new animal species. While for confusing concepts, the problem usually lies with biased concept representation in web image-text data. For example: online images of “Resupply Base” mostly only consist of exterior views of the architecture without the interior details, which may cause the models to confuse it with a “Wholesale Store” that shares some interior features.
To help LMMs recognize and reason about novel and confusing concepts more robustly and efficiently, we propose CoDA, a Contrastive Visual Data Augmentation technique. For each target concept, CoDA first identifies a “confusable concept” that the LMMs finds most similar to the target. Then, it extracts contrastive textual and visual features of the target concept with respect to the confusable concept. The extracted features go through a filtering process based on discriminability and generability to make sure that: 1). The features are possessed by the target concept but not the confusable concept; and 2). The feature can be reliably generated by the text-to-image generative model and recognized by the LMMs. Afterwards, the features are passed to the text-to-image generative model to produce augmented visual instances of the target concept. To make sure that the features are indeed generated and recognizable by the LMMs, CoDA again uses the LMMs’ zero-shot inference to rank and filter the augmented images. Finally, the resulting augmented images can be used to update the LMMs via low-rank adaptation, basic fine-tuning, in-context learning, or any other method of choice.
@@
Comprehensive experiments with LLaVA-NeXT on the 3 datasets show CoDA performs surprisingly well in teaching LMMs novel and confusing concepts, significantly improving data efficiency compared to existing methods. In additional experiments, we show that CoDA is also able to improve novel concept recognition for traditional classifiers like ViT and proprietary LMMs such as GPT4o-mini. Finally, ablation experiments show that CoDA can be significantly improved by simply replacing its off-the-shelf components such as the text-to-image generation model with superior versions of similar models.
@@
  • NovelSpecies, a new benchmark dataset of novel animal species discovered in recent years, providing an ideal benchmark for novel concept recognition. NovelSpecies currently consists of 2240 annotated images and will continue to be updated with future discoveries.
@@ 2 Related Works
Few-shot image recognition is a long-standing problem in the vision community. Early works in this area focused on improving traditional image classifiers on classifying existing concepts (Vinyals et al., 2016; Finn et al., 2017; Nichol, 2018; Dhillon et al., 2019; Tian et al., 2020; Bhagat et al., 2023; Afrasiyabi et al., 2022). On the other hand, while recent advancements in the training of vision language models (VLMs) and large multimodal models (LMMs) (OpenAI, 2023; Google, 2023; Hurst et al., 2024; Liu et al., 2023b, a, 2024c) have shown great promise and extensibility, they still severely lag behind traditional models in image classification, especially for low-resource, novel, and confusing concepts (Zhang et al., 2024; Cooper et al., 2024; Wu et al., 2023b; Yang et al., 2024; Ha et al., 2025).
While commonly used text-side VLM data augmentation strategies (Yuksekgonul et al., 2022; Yang et al., 2023; Liu et al., 2024e; Sharifzadeh et al., 2024) have little effect on this issue, a more promising technique to solve this is through visual data augmentation. This includes basic visual manipulations such as cropping, flipping, and rotation (Yang et al., 2022; Kumar et al., 2024); and more advanced model-based augmentation such as style transfer (Zheng et al., 2019; Chun & Park, 2021) and image mixing (Uddin et al., 2020; Xie et al., 2021; Hao et al., 2023). More recently, with the rise of controllable and promptable visual generative models, knowledge and feature editing-based augmentation methods (Liu et al., 2022; Wu et al., 2023a; Jin et al., 2024) have gained in popularity. Such methods generally focus on using multimodal data and general knowledge bases to guide image-editing models in creating augmented visual data based on existing images.
@@ 3 Methods
As shown in Figure 2, CoDA consists of 4 major steps including contrastive textual and visual feature extraction, feature filtering, feature-controlled image generation, and augmented image filtering. Together these steps ensure CoDA reliably generates informative and high-quality augmented images that help LMMs recognize novel and confusing concepts.
@@ 3.1 Feature Extraction
In our exploratory experiments, we find that significant mis-recognition errors occur on low-resource or commonly mis-represented concepts in vision-language instruction fine-tuning and multimodal pre-training datasets, which the LMMs are trained on. For example, the LLaVA 1.6 (34B) model (Liu et al., 2024c), mainly tuned on LAION-GPT-4V(LAION, 2024) and ShareGPT-4V (Chen et al., 2023) datasets, has a strong tendency to mis-recognize interior images of “Resupply Base” as “Wholesale Store” (Figure 1). Unsurprisingly, we find that all related references of “Resupply Base” across the 3 instruction-tuning datasets only depict exterior views of the concept rather than interior views. While the concept itself is not a low-resource concept in existing text corpora, it is severely low-resource and also commonly mis-represented in vision-language instruction fine-tuning datasets.
@@
While textual feature extraction generally works well for pre-existing and non-hyper-domain-specific concepts that are prevalent in textual data sources, it tends to fail when either of the conditions are not met. For example, a large language model with a knowledge cutoff prior to June 2023 would not be able to provide meaningful features regarding the Apple Vision Pro device announced in July, or the new animal species “Clouded Tiger Cat (L. pardinoides)” first described by scientists in April 2024 (Figure 1). In addition to this weakness, LLM-based textual feature extraction is also unreliable when asked to provide detailed information regarding hyper-domain-specific concepts like the “Mazda MX-5 Miata RF” or the “Lear’s Macaw (Anodorhynchus Leari)”. In practice, we observe that for novel and hyper-domain-specific concepts, most of the LLM extracted textual features end up being filtered out by our automatic feature filtering module.
@@ 3.2 Feature Filtering
    Discriminability ($D(f,\mathcal{C}_{T},\mathcal{C}_{M})$) : measures whether a feature $f$ indeed differentiates the target class $\mathcal{C_{\mathcal{T}}}$ from the misidentified concept $\mathcal{C_{\mathcal{M}}}$ (check whether $f$ is a valid feature of $\mathcal{C_{\mathcal{T}}}$ but not $\mathcal{C_{\mathcal{M}}}$) .
@@
Here we use an equal number of images of the target and misidentified concepts. A score below 0.5 indicates that the feature is more likely to be associated with the misidentified class rather than the target class. To ensure that selected features are more strongly associated with the target class, we filter out all features with Discriminability below 0.6. This method avoids the CLIP score bias against smaller features by only comparing feature association with the two classes and not relying on the absolute CLIP score.
Generability is calculated for all features that pass the Discriminability threshold. We prompt the T2I generative model $g$ to generate synthetic images of the target concept that contains the feature $f$, and then compare the average CLIP similarity between $f$ and those generated images against the average CLIP similarity between $f$ and $i_{\mathcal{C}_{M}}^{\text{real}}$:
$G(f,\mathcal{C}_{T},\mathcal{C}_{M},g)=\sum_{i\in I}\frac{\text{CLIP}(f,i_{\mathcal{C}_{T}}^{\text{synthetic}})}{\text{CLIP}(f,i_{\mathcal{C}_{T}}^{\text{synthetic}})+\text{CLIP}(f,i_{\mathcal{C}_{M}}^{\text{real}})}$
@@ 3.3 Image Generation and Verification
To ensure final images for augmentation contain our extracted and filtered target concept features, we propose a simple automatic verification metric that checks whether desired features are recognized in the augmented images by the LMM we want to update: Given the vanilla LMM $\mathcal{M}$, a set of features $\mathcal{F}$, and an augmented image $i^{\text{synthetic}}$, the feature satisfaction rate $S(i^{\text{synthetic}},F,M)$ for each augmented image:
@@ 3.4 Human Evaluation
Results in Table 1 show human annotators overwhelmingly agree that the final extracted features belong to the target concept (92%) but not the misidentified concept (14%). The augmented synthetic images of the target concept also likely contains the desired features (83%), though as expected, there is some information loss between the text-to-image generation step. In addition, the three independent annotators generally agreed in their response (>0.8 IAA).
@@ 4 NovelSpecies Dataset
To create NovelSpecies, we start by collecting the list of species first described in each year by Wikidata (Wikidata, 2024). Then, to make sure we can curate a visual benchmark of novel species, we manually annotate and filter out extinct species and species with too few publicly available images. After filtering, we end up with a dataset of 64 new species, each consisting of 35 human-verified image instances, thus a total of 2240 images. The images are split into training, validation, and test sets. For each species, there are 5 training images, 15 validation images, and 15 test images. This data split is consistent with our goal of creating a benchmark dataset for novel concept recognition, where the maximum number of training instances for a completely unseen concept can range from 1 to 5.
@@ 5.1 Datasets and Baselines
In addition to these 3 baselines, we also include ablations of CoDA with non-contrastive textual and visual features, i.e. w/o contrastive guidance from confusable concepts (3.1) nor discriminability-based feature filtering (3.2).
@@ 5.2 Main Experiment
Experiment results across the 3 datasets show that CoDA consistently outperforms existing traditional and feature-based data augmentation methods in the Fixed Real Data setting. When augmenting the training set with just a single synthetic image, CoDA is able to achieve 11.8% (NovelSpecies), 10.0% (SUN), and 17.8% (iNat) absolute gains in accuracy compared to using all real images. It further outperforms the best existing baseline augmentation methods by 5-12% absolute gains. We also observe that the ablated performance of CoDA (w/o contrastive) is still significantly above traditional and image-editing-based augmentation baselines while being almost consistently below CoDA’s performance. This shows the benefits of text-to-image generative augmentation methods compared to existing methods, as well as the benefits of fine-grained textual features during inference. This also highlights the need for contrastive feature selection and discriminability-based feature filtering. We find that increasing the number of augmented synthetic images does not necessarily improve updated model performance; this may be attributed to the fact that all generated images are ranked and selected from the same pool, with the first image being of the highest quality. Finally, the largest improvement over existing baselines can be seen in NovelSpecies, where CoDA methods involving visual features achieve the highest performance. This makes sense as the visual feature extraction method is designed to be robust to novel concepts with little textual documentation.
@@ 5.3 Additional Experiments
As explained in Sec.3, off-the-shelf model components used in CoDA can be easily swapped for superior versions of similar models to improve performance. To demonstrate this, we replace the open-weight Stable Diffusion 3.5 Large Turbo model (Stability AI, 2024) with the SOTA proprietary Recraft V3 Model (Recraft.AI, 2024) and run the same LLaVA-updating experiments as in Table 2. Here we note that Recraft V3 has better instruction-following ability as well as better image generation quality compared to Stable Diffusion 3.5 Large Turbo. More details on these differences can be found in Sec.6. Our experiment results in Table 3 show a significant performance boost when LoRA fine-tuning LLaVA with Recraft V3 produced synthetic images compared to fine-tuning on all-real data (28.7%) and also compared to fine-tuning on Stable Diffusion 3.5 Large Turbo produced synthetic data (7.9%). This demonstrates the potential increase of CoDA’s effectiveness along with improvements in Text-to-Image generative models. We believe it is also possible to achieve similar improvements by replacing the LLM/VLM components of CoDA with superior models in the future.
@@
While proprietary LMMs like GPT4o-mini (Hurst et al., 2024) tend to have relatively strong 0-shot performance on existing datasets such as SUN and iNaturalist, their performance significantly degrades on NovelSpecies due to having never encountered the novel concepts. To test whether CoDA can effectively improve novel concept recognition performance for such proprietary LMMs, we fine-tune the gpt-4o-mini-2024-07-18 model using CoDA and relevant augmentation baselines. Results in Table 3 demonstrate a significant performance gain (9.5%) for GPT4o-mini after being fine-tuned on CoDA augmented synthetic images. While this improvement is not as significant compared to the LLaVA-1.6 model (20.3%), it is due to GPT4o-mini’s better base performance.
@@
In addition to evaluating CoDA on LMMs which take image-text input and produce text output, we also test whether it can help traditional image classifiers recognize novel concepts. We run the widely-used ViT classifier (Alexey, 2020) on NovelSpecies with CoDA and other augmentation baselines. Results in Table 3 show that CoDA is able to achieve a consistent performance gain over existing baselines for ViT-base (9.1% for single-shot augmentation). The ViT classifier provides stronger base performance compared to general VLMs, thus offering less room for improvement. However, we note here that our main focus on improving LMMs instead of such traditional classifiers stems from LMMs’ superior extensibility and generalizability to other related tasks such as recognition-based reasoning and explanation.
6 Discussions
Figure 3: Qualitative Comparison of CoDA and baseline visual data augmentation methods. Phyllobates Samperi and Tail-Spot Wrasse are example concepts from the NovelSpecies dataset. All CoDA images are generated using contrastive textual + visual features.
In Figure 3, we compare example synthetic images generated by CoDA and baseline visual data augmentation methods including Cropping and ARMADA (Jin et al., 2024). While providing localized feature emphasis, Cropping often results in the loss of crucial visual details necessary for concept identification. For instance, for Phyllobates Samperi, cropping occludes the black spots on the frog’s skin, an essential distinguishing feature. Without such essential distinguishing features, the cropped images provide less helpful learning signals compared to other methods.
Unlike Cropping, ARMADA successfully retains some structural features of the target concept, using WikiData text features to guide its image-editing backbone (Brooks et al., 2023). However, this setup also induces two significant issues: (1) Leveraging only existing textual features present in WikiData leads to an incomplete feature set, especially for novel concepts. This is apparent in images generated for Phyllobates Samperi, where generated images contained ”black spots” but failed to specify their location on the legs of the frog instead of the body. In addition, the model completely failed to generate Phyllobates Samperi’s iconic ”orange stripes” due to the feature not being recorded in WikiData. (2) Image-editing models are not as strong in depicting precise details compared to text-to-image generative models. In the case of the Tail-Spot Wrasse, the ARMADA generated image fails to accurately depict ”vibrant horizontal stripes along the body,” leading to a visually inconsistent and less biologically accurate representation.
In contrast to existing baselines, Figure 3 shows that CoDA is much better at generating high quality synthetic images of the target novel concept that depict accurate and realistic details. Both versions of CoDA using different backbone models (Esser et al., 2024; Recraft.AI, 2024) are able to produce significantly more realistic images compared to the two baselines, maintaining general biological consistency. However, we should note that CoDA’s performance is inherently bounded by the instruction-following ability of its image generation backbone model, more specifically the ability to accurately generate multiple feature details in a single image. For example, while CoDA-Recraft-V3 is able to accurately generate all three extracted features including ”prominent orange stripes”, ”dark shiny coloration”, and ”black spots on legs”; CoDA-SD-3.5 is only able to generate the first two features while failing to capture ”black spots on legs”. With such limitations in mind, we give CoDA an extremely modularized design. This allows each pre-trained model component in CoDA to be easily replaced for newer and stronger versions of similar models, including more perceptive VLMs and T2V generative models with stronger instruction-following ability and higher generation quality.